Christian Golbach stated that every odd number greater than 7 can be written as the sum of three prime numbers. rewrite each of the following as the sum of three prime numbers: 77, 115, and 249.
Answer by richard1234(7193)

77 = 5+5+67
115 = 3+5+107
249 = 7+31+211

In fact, the weak Goldbach Conjecture states that every odd number greater than 7 can be written as the sum of three "odd" prime numbers (which means that 2 would be excluded). However this, as well as the original Goldbach Conjecture, remain unsolved.